# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

### v2

@@ III-B2 DPLib AC OPF Solver
The DPLib AC OPF solver implements a distributed ADMM-based solver in MATLAB for any power system preprocessed using the DPLib partitioning tool. It enables large-scale distributed analysis, where each region independently solves its local AC OPF while coordinating inter-regional tie-lines. The solver developed in this work is publicly available at GitHub, including the MATLAB implementation file named DistACOPF.m. To run the code, the user must have MATLAB R2020a or later, with MATPOWER and IPOPT properly configured. Upon launch, the code prompts for the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After loading the partitioned data, the code augments each region’s branch matrix with its tie-lines and initializes region-specific variables (e.g., bus counts, initial voltage values). The main ADMM loop iteratively solves each region’s local AC OPF via the helper function Acopf, based on (2), updating voltage magnitudes and angles. Dual variables are adjusted using consistency gradients to enforce agreement across tie-lines. Convergence is declared if the worst primal residual falls below the threshold. The optimality gap is computed per iteration if a centralized cost is provided. Final plots show convergence trends, and status messages are printed throughout. The overall process flow of the DPLib AC OPF solver is presented in algorithm 2.
@@ IV DPLib: Benchmark Test Cases
Distributed versions of many widely used MATPOWER benchmark systems are generated. Each test case is processed using the DPLib graph-based partitioning tool and tested using the DPLib ADMM-based DC and AC OPF solvers described earlier. Across all systems, both solvers successfully converge to feasible solutions. All partitioned cases and solver outputs are publicly available as part of the DPLib at GitHub. In the following subsections, we explain some of the distributed test cases generated and tested as benchmarks in DPLib.
In addition to evaluating convergence using the worst-case primal residual and the optimality gap, further validation tests are performed on the generated multi-region systems to verify their consistency with physical laws and operational constraints of the power grid. Specifically, we verified that upon the convergence of the ADMM iterations, the total power generation across all regions closely matches the total system demand, accounting for minor discrepancies due to modeling approximations and convergence thresholds. Furthermore, we confirmed that active power flows across each inter-regional tie-line are consistent from both sides of the interface. All tested datasets successfully passed these validation checks, further demonstrating the correctness and reliability of the generated partitioned cases.
To facilitate evaluation of the benchmark test cases using OPF solvers, all primal variables were initialized with cold-start conditions: voltage angles were set to zero and voltage magnitudes to one per unit. The dual variables, corresponding to Lagrange multipliers associated with consistency constraints, were initialized to one. Penalty parameters $\rho$ in the ADMM-based solvers were selected case by case to ensure convergence.
IV-A Case: pglib_opf_case200_tamu
The pglib_opf_case200_activ system is a synthetic 200-bus benchmark from the IEEE PES Power Grid Library for optimal power flow studies. Geographically situated in the central region of Illinois, this case models a 230/115 kV transmission network designed to serve a load distribution that mimics the population profile of its geographic footprint. The transmission system was generated following the methods outlined in [19], ensuring statistical similarity to real-world grids without modeling any actual transmission lines. Developed as part of the U.S. ARPA-E GRID DATA research project, the system contains no critical energy infrastructure information and was created using PowerWorld Simulator, then converted to MATPOWER format using version 6. The dataset is publicly available through PGLib and licensed under the Creative Commons Attribution 4.0 International License.
The pglib_opf_case200_tamu test case was partitioned using our graph-based spectral clustering method, which ensured full bus inclusion and topological connectivity. The best clustering configuration identified by the algorithm consisted of three electrically coherent regions, interconnected through 10 tie-lines, within 10 iterations. During the partitioning, the reference bus, bus 189, was assigned to Region 2 (mpc_regionR2). No buses were omitted, and the total number of buses across all regions matches the original MATPOWER dataset, with 200 buses, 245 branches, and 49 generators. The consistency of branch count, generator count, and overall system topology confirms the fidelity of the partitioning process. The final multi-region topology has been visualized and is shown in Fig. 2. The corresponding partitioned files are saved as pglib_opf_case200_tamu_3regions.mat and pglib_opf_case200_tamu_3regions.m. Both distributed DC OPF and AC OPF solvers have been executed on this partitioned data, and their convergence trends are illustrated through the worst primal residuals presented in Fig. 3 with $\rho=1000$.
Figure 2: Topology of pglib_opf_case200_tamu partitioned into three regions with 10 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 3: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case200_tamu.
IV-B Case: pglib_opf_case300_ieee
The pglib_opf_case300_ieee system is a widely used benchmark from the IEEE PES Power Grid Library for optimal power flow studies. This test case represents the IEEE 300-bus system, which was initially provided in the IEEE Common Data Format and converted into MATPOWER format on September 20, 2004, using the cdf2matp utility. The dataset was sourced from the University of Washington’s power systems test case archive. The system corresponds to typical operational conditions, with a base power of 100 MVA and includes detailed modeling of generation, loads, and transmission elements.
Using our graph-based spectral clustering approach, we have partitioned the pglib_opf_case300_ieee system into three electrically coherent regions. Before clustering, the original bus numbering in the MATPOWER file was reordered to ensure consistent indexing. The graph topology of the network was verified to be fully connected, allowing for a valid partitioning process. The algorithm identified the best clustering configuration in 10 iterations, resulting in 3 regions interconnected through 12 tie-lines. The system’s slack bus, originally defined at bus 257, was successfully assigned to Region 1 (denoted as mpc_regionR1). After partitioning, the integrity of the dataset was confirmed: all 300 buses, 411 branches, and 69 generators were fully included in the regional breakdown, with no data omitted. The final partitioned system has been saved as both a .mat file and a script-based .m file named pglib_opf_case300_ieee_3regions. A visual representation of the multi-regional topology is shown in Fig. 4. The distributed DC OPF and AC OPF solvers were successfully executed on this dataset, and the convergence behavior is demonstrated through the worst primal residual plots shown in Figs. 5(a) and 5(b) with $\rho=1000$.
Figure 4: Topology of pglib_opf_case300_ieee partitioned into three regions with 12 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 5: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case300_ieee.
IV-C Case: pglib_opf_case500_tamu
The pglib_opf_case500_tamu case is a synthetic 500-bus power system in the IEEE PES Power Grid Library for OPF studies. This case was developed as part of the ARPA-E GRID DATA research initiative and is geographically based in the northwestern region of South Carolina. It incorporates a transmission network operating at 345 kV and 138 kV voltage levels, constructed to supply power to a synthetic load distribution that mirrors the population density of the modeled region. The system does not represent any real grid infrastructure but is created using statistical modeling techniques described in [19], ensuring structural similarities to real transmission systems while maintaining confidentiality.
Using our graph-based spectral clustering framework, the pglib_opf_case500_tamu benchmark system was partitioned into five electrically coherent regions. Before clustering, the bus numbers from the original MATPOWER file were analyzed to preserve consistent indexing. The graph representation of the system was confirmed to be fully connected, ensuring the validity of the partitioning procedure. The spectral clustering algorithm converged in 10 iterations and identified an optimal structure of 5 regions interconnected through 16 tie-lines. The reference bus, originally defined as bus 17, was successfully assigned to Region 2 (denoted as mpc_regionR2). Post-partitioning checks confirmed that all 500 buses, 597 branches, and 90 generators were preserved without omission or duplication. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case500_tamu_5regions. A visual illustration of the multi-regional layout is presented in Fig. 6. Distributed DC OPF and AC OPF solvers were successfully executed for this case, and the convergence profiles are illustrated by the worst primal residual curves in Figs. 7(a) and 7(b) with $\rho=1000$.
Figure 6: Topology of pglib_opf_case500_tamu partitioned into five regions with 16 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 7: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case500_tamu.
IV-D Case: pglib_opf_case1354_pegase
The pglib_opf_case1354_pegase dataset is a detailed test case that emulates the size and complexity of the European high-voltage transmission network. It includes 1,354 buses, 260 generators, and 1,991 branches operating at 220 and 380 kV voltage levels. While the data is fictitious and does not reflect real-world infrastructure, it is highly suitable for validating power system optimization tools and techniques. This benchmark was developed as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project. It incorporates unique modeling features such as asymmetric shunt elements and generator aggregations.
The pglib_opf_case1354_pegase benchmark system was partitioned into eight electrically coherent regions using our graph-based spectral clustering method. The original bus indices from the MATPOWER file were reordered to ensure consistency during partitioning. The graph representation of the power network was verified to be fully connected, ensuring that a valid and complete partitioning could be achieved. The clustering algorithm successfully identified an optimal partitioning configuration within 10 iterations, resulting in 8 regions interconnected through 62 tie-lines. The system’s reference bus, which is located at bus 640 in the original data, was assigned to Region 1 (denoted as mpc_regionR1). The integrity of the data was verified after partitioning: all 1354 buses, 1991 branches, and 260 generators were accurately distributed across the regions with no missing components. The partitioned data files were saved in both .mat and .m formats under the name pglib_opf_case1354_pegase_8regions. A graphical representation of the resulting multi-regional topology is provided in Fig. 8. Both distributed DC OPF and AC OPF solvers were applied to this case study, with convergence behavior illustrated via worst primal residual plots in Figs. 9(a) and 9(b) with $\rho=1000$.
Figure 8: Topology of pglib_opf_case1354_pegase partitioned into eight regions with 62 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 9: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case1354_pegase.
IV-E Case: pglib_opf_case2383wp_k
The pglib_opf_case2383wp_k dataset models the Polish national transmission grid under peak load conditions during the winter of 1999–2000. It includes a comprehensive representation of 400, 220, and 110 kV networks and originates from a reduced version of the broader 7500+ bus European UCTE system. To simplify the model while preserving essential characteristics, tie lines connecting to neighboring countries were replaced by artificial load or generator buses. Generator aggregation and cost approximations were applied to reflect dispatchability in the Polish energy market.
Using the proposed graph-based spectral clustering method, the pglib_opf_case2383wp_k benchmark system was partitioned into twelve electrically meaningful regions. The underlying graph structure of the power system was verified to be fully connected, confirming the feasibility of multi-regional partitioning. The algorithm determined the optimal clustering within 10 iterations, producing 12 regions linked by 99 inter-regional tie-lines. The system’s slack bus, originally defined at bus 18 in the MATPOWER case, was assigned to Region 4 (referred to as mpc_regionR4).
Following partitioning, a full consistency check verified that all 2383 buses, 2896 branches, and 327 generators were preserved and correctly distributed across the regional datasets. No components were omitted during the process. The results were stored as both a .mat file and an executable .m script, saved under the name pglib_opf_case2383wp_k_12regions. The partitioned network topology is visually represented in Fig. 10. The distributed DC OPF and AC OPF solvers are implemented and their convergence behavior is shown through the worst primal residual plots in Figs. 11(a) and 11(b) with $\rho=2000$.
Figure 10: Topology of pglib_opf_case2383wp_k partitioned into twelve regions with 99 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 11: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case2383wp_k.
IV-F Case: pglib_opf_case3012wp_k
The pglib_opf_case3012wp_k dataset provides a detailed model of the Polish transmission grid during the winter 2007–2008 evening peak demand period. It represents the operational state of the 400, 220, and 110 kV networks and is a high-fidelity snapshot of the Polish system during that time. As part of the Power Grid Library benchmark suite, it is widely used for testing and validating optimization algorithms in realistic, large-scale scenarios.
Using the proposed graph-based spectral clustering framework, the pglib_opf_case3012wp_k benchmark system was partitioned into twelve electrically coherent regions. Before partitioning, the original MATPOWER bus numbers were reordered to enable consistent indexing throughout the graph representation. The full graph connectivity was verified to ensure that all buses could be properly included in the clustering process. Within 10 iterations, the partitioning algorithm identified an optimal configuration that divides the system into 12 regions connected by 100 inter-regional tie-lines. The slack bus, located initially at bus 37 in the MATPOWER data, was assigned to Region 12 (denoted as mpc_regionR12).
Following the partitioning process, a verification step confirmed that the distributed representation included all 3012 buses, 3572 branches, and 502 generators. These components were preserved without omission or duplication. The resulting data was saved both as a .mat file and as a script-based .m file under the name pglib_opf_case3012wp_k_12regions. A visual overview of the final partitioned topology is shown in Fig. 12. Using this distributed dataset, the DC OPF and AC OPF solvers were executed successfully, with their convergence behavior illustrated by the worst primal residual plots shown in Figs. 13(a) and 13(b) with $\rho=2000$.
Figure 12: Topology of pglib_opf_case3012wp_k partitioned into twelve regions with 100 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 13: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case3012wp_k.
IV-G Case: pglib_opf_case4661_sdet
The pglib_opf_case4661_sdet dataset is a large-scale synthetic power system model with 4,661 buses. It was created using the Sustainable Data Evolution Technology tool and is publicly available through the eGRID Data portal. This model is part of the PGLib OPF v19.05 benchmark suite and is provided in the public domain to support research in scalable and realistic power system analysis.
The pglib_opf_case4661_sdet benchmark system was partitioned into fourteen electrically coherent regions using the proposed graph-based spectral clustering methodology. Before partitioning, the original bus numbering in the MATPOWER data was reordered to enable consistent and systematic region assignment. The underlying graph representation of the network was verified to be fully connected, ensuring that all components were eligible for inclusion in the partitioning process. After 10 iterations, the algorithm identified an optimal clustering configuration with 14 regions and 284 inter-regional tie-lines. The reference (slack) bus, originally located at bus 2625 in the MATPOWER data, was assigned to Region 3 (denoted as mpc_regionR3).
A validation step was performed after partitioning to confirm the structural integrity of the distributed dataset. It verified that all 4661 buses, 5997 branches, and 1176 generators from the original MATPOWER case were successfully preserved and assigned across the generated regional files. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case4661_sdet_14regions. A visual summary of the multi-regional topology is presented in Fig. 14. Subsequently, the distributed DC OPF and AC OPF solvers were executed on the partitioned system, and their convergence characteristics are shown through the worst primal residual plots in Figs. 15(a) and 15(b) with $\rho=4000$.
Figure 14: Topology of pglib_opf_case4661_sdet partitioned into fourteen regions with 284 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 15: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case4661_sdet.
IV-H Case: pglib_opf_case6468_rte
The pglib_opf_case6468_rte [20] dataset provides a realistic snapshot of the French high voltage and very high voltage transmission network in 2013. It contains 6,468 buses, 1,295 generators, and 9,000 branches, offering a highly detailed and complex test case for power system optimization. While the data are based on real grid topologies, they are intended solely for academic and research purposes and should not be used for operational or planning decisions related to the French or European power grids. This dataset is part of the IEEE PES Power Grid Library (PGLib-OPF v19.05) and was developed by Réseau de Transport d’Électricité (RTE), the French transmission system operator. It supports validating optimization algorithms, mathematical modeling techniques, and simulation tools in large-scale power system analysis.
The pglib_opf_case6468_rte benchmark was successfully partitioned into electrically coherent regions using our spectral clustering-based partitioning algorithm. This method ensured that the network’s graph structure remained fully connected. The best clustering outcome was identified within 10 iterations, resulting in 16 regions interconnected by 204 inter-regional tie-lines. Bus renumbering was applied to the original MATPOWER case to enable consistent indexing across the partitioning workflow. The system’s slack bus, labeled as bus 4733, was assigned to Region 6 (mpc_regionR6).
All buses, branches, and generators from the original MATPOWER dataset were accounted for in the partitioned output, confirming the completeness and validity of the partitioning. Specifically, the resulting distributed dataset includes all 6468 buses, 9000 branches, and 1295 generators from the original case. The partitioned version has been saved in both .mat and script-based .m formats under the name pglib_opf_case6468_rte_16regions. A visualization of the region-level partitioned topology is presented in Fig. 16. To validate the partition, we applied the distributed DC OPF and AC OPF solvers, which successfully converged. Their worst primal residual plots over ADMM iterations are shown in Fig. 17 with $\rho=6000$.
Figure 16: Topology of pglib_opf_case6468_rte partitioned into 16 regions with 204 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 17: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case6468_rte.
IV-I Case: pglib_opf_case9241_pegase
The pglib_opf_case9241_pegase dataset provides a highly detailed representation of the European high voltage transmission network. It includes 9,241 buses, 1,445 generators, and 16,049 branches operating across multiple voltage levels, including 750, 400, 380, 330, 220, 154, 150, 120, and 110 kV. Although the dataset is fictitious and does not reflect real-world data, it is valuable for validating optimization algorithms and simulation tools in large-scale power system research. The data were released as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project [21].
The pglib_opf_case9241_pegase benchmark represents one of the largest systems in the MATPOWER repository and was successfully partitioned using our graph-based spectral clustering methodology. The partitioning process ensured complete inclusion of all 9241 buses and confirmed the full connectivity of the underlying power grid graph. The best clustering configuration was determined within 10 iterations, resulting in 18 regions connected by 197 inter-regional tie-lines. The system’s reference bus, numbered 4231, was assigned to Region 17 (mpc_regionR17). The reorganization of the data preserves the integrity of the original MATPOWER structure, with a total of 16049 branches and 1445 generators fully accounted for across the distributed regions. The partitioned results were saved in both .mat and .m formats under the name pglib_opf_case9241_pegase_18regions. A visual representation of the multi-region topology is provided in Fig. 18. This dataset has been validated by executing the distributed DC OPF and AC OPF solvers, with the convergence of the worst primal residuals depicted in Fig. 19 with $\rho=8000$.
Figure 18: Topology of pglib_opf_case9241_pegase partitioned into 18 regions with 197 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 19: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case9241_pegase.

### v3

@@ III-B2 DPLib AC OPF Solver
The DPLib AC OPF solver provides a MATLAB implementation of a distributed ADMM-based AC OPF algorithm that operates on any power system previously partitioned using the DPLib framework. The solver is implemented in the publicly available file DistACOPF.m, accessible through the GitHub repository at GitHub. The code requires MATLAB R2020a or later, along with a properly configured installation of MATPOWER and IPOPT.
The function begins by requesting the partitioned .mat file and the user-defined ADMM parameters, including the penalty value $\rho$, the convergence tolerance, the maximum number of iterations, the $\rho_{\mathrm{AC}}^{\max}$, and optionally the centralized AC OPF cost for computing the optimality gap. After these inputs are provided, the solver loads the partitioned data and augments the branch matrices of each region with their associated tie-lines. It then initializes regional variables describing bus counts, voltage magnitudes and angles, and other quantities needed to enable the local AC OPF solves and the coordination steps.
The main ADMM loop is implemented within the DistACOPF.m function. In each iteration, every region independently solves its local nonlinear AC OPF using the helper function acopf, which is based on the formulation in equation (4). The updated voltage magnitudes and phase angles from each region are then exchanged virtually through the consistency constraints applied on the tie-lines. The dual variables are modified at each step using the mismatch in the angle and magnitude variables across the interfaces, ensuring that neighboring regions gradually reach agreement. The loop continues until the maximum primal residual across all tie-lines drops below the specified tolerance. If a centralized AC OPF cost is provided, the function also evaluates the optimality gap at every iteration to quantify the progress toward the global optimum.
Once the iterations terminate, the solver prints summary information and generates plots that illustrate the convergence behavior of the primal residual and, when applicable, the optimality gap. The complete algorithmic flow implemented in DistACOPF.m is summarized in Algorithm 2.
@@ IV DPLib: Benchmark Test Cases
TABLE I: Partitioning summary with centralized DC/AC OPF costs.
System Size Partition Penalty Param. Centralized OPF Cost
Case Buses Branches Gens Regions Tie-lines Slack $\rho$ $\rho_{\mathrm{DC}}^{\max}$ $\rho_{\mathrm{AC}}^{\max}$ DC AC
pglib_opf_case200_tamu $200$ $245$ $49$ 33 $10$ R2 11 1e81108 1e81108 $34,653$ $27,558$
pglib_opf_case300_ieee $300$ $411$ $69$ 33 $12$ R1 11 1e81108 1e81108 $517,310$ $565,220$
pglib_opf_case500_tamu $500$ $597$ $90$ 55 $16$ R2 11 1e81108 1e81108 $91,841$ $72,578$
pglib_opf_case1354_pegase $1,354$ $1,991$ $260$ 88 $62$ R1 11 1e81108 1e81108 $1,218,028$ $1,258,844$
pglib_opf_case2869_pegase $2,869$ $4,582$ $510$ $12$ $95$ R6 11 1e51105 1e81108 $2,386,170$ $2,462,790$
pglib_opf_case4661_sdet $4,661$ $5,997$ $1,176$ $14$ $284$ R3 11 1e61106 1e61106 $2,217,130$ $2,255,540$
pglib_opf_case9241_pegase $9,241$ $16,049$ $1,445$ $18$ $197$ R17 11 1e61106 1e51105 $6,044,174$ $6,243,090$
This section presents a detailed description and verification of several DPLib benchmark systems and their corresponding multi-region topologies. All test cases originate from the PGLib-OPF library and are partitioned using the spectral clustering procedure of Algorithm 1 based on the unweighted Laplacian matrix. Before partitioning, every system undergoes bus renumbering when required, connectivity verification, slack-region assignment, and integrity checks to ensure that all buses, branches, and generators are preserved. The resulting number of regions, tie-lines, and solver parameters are summarized in Table I. All region-labeled MATPOWER files, verification codes, and the topology and verification figures included here are available in the DPLib GitHub repository.
To make the library easy to use and to avoid case-by-case tuning, the ADMM penalty parameters are initialized to $\rho^{0}=1$ for all benchmarks in both the distributed DC-OPF and AC-OPF solvers, after which $\rho$ is adjusted automatically via the adaptive residual-balancing rule where $\tau_{\mathrm{incr}}=2,\tau_{\mathrm{decr}}=0.5$. In the distributed DC-OPF solver, the adaptive update parameters are set internally to
$\mu=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),$ (8)
where $n_{\mathrm{reg}}$ denotes the number of regions. For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and voltage-magnitude channels using
$\mu_{\theta}=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),\;\mu_{v}=\max\!\left(1,\ \frac{10}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right).$ (9)
These choices scale the residual-balance thresholds with the partition granularity while keeping fixed and interpretable increase/decrease factors, allowing the adaptive-$\rho$ mechanism to operate robustly across different network sizes and region counts with minimal user intervention.
The verification procedure is implemented in a MATLAB script that processes partitioned cases in a unified workflow. For each system, the script first loads the corresponding MATPOWER case and calls the centralized solvers run_dcopf_centralized and run_acopf_centralized. These functions compute reference DC OPF and AC OPF objective values and return solver status information. The centralized costs serve as benchmarking values to compute the optimality gap of the distributed solutions if the centralized DC OPF is feasible. The distributed DC OPF solver is run using the function DistDCOPF with prescribed tolerance and maximum iteration count. Similarly, the distributed AC OPF solver DistACOPF is called following the structure of Algorithm 2. For each distributed run, the script tracks the worst primal residual across tie-lines and the optimality gap with respect to the centralized objective value. Table I reports placeholder centralized DC and AC objective values for all systems. In the numerical experiments, for all cases where the centralized solvers return feasible solutions, the distributed solvers drive the primal residual down to the threshold and achieve an optimality gap below one percent, indicating that the distributed solutions are both feasible and near-optimal.
Case pglib_opf_case200_tamu
The system pglib_opf_case200_tamu is a 200-bus synthetic model representing a 230/115 kV network in Illinois. It was generated as part of the ARPA-E GRID DATA project and is designed to capture realistic transmission characteristics and operational patterns. The system is partitioned into three regions connected through ten tie-lines. The resulting multi-region topology is shown in Fig. 3.
Figure 3: Three-region topology for case200_tamu.
Figures 4 and 5 report the DC and AC primal residuals and optimality gaps for pglib_opf_case200_tamu.
Figure 4: primal residual (left) and optimality gap (right)
Figure 5: primal residual (left) and optimality gap (right)
Case pglib_opf_case300_ieee
The system pglib_opf_case300_ieee corresponds to the well-known IEEE 300-bus test case. Originally formulated in the IEEE Common Data Format and subsequently converted to MATPOWER, it has long served as a standard benchmark for power system optimization studies. The system is divided into three regions, with interconnections distributed across twelve tie-lines. The extracted topology is shown in Fig. 6.
Figure 6: Three-region topology for case300_ieee.
Figures 7 and 8 show the DC and AC verification results for pglib_opf_case300_ieee.
Figure 7: primal residual (left) and optimality gap (right)
Figure 8: primal residual (left) and optimality gap (right)
Case pglib_opf_case500_tamu
The system pglib_opf_case500_tamu is a 500-bus synthetic model of the South Carolina transmission grid, also produced under the ARPA-E GRID DATA initiative. It contains detailed line, transformer, and generator data reflecting real-world system behavior. The network is partitioned into five regions connected by sixteen tie-lines. The corresponding topology appears in Fig. 9.
Figure 9: Five-region topology for case500_tamu.
For pglib_opf_case500_tamu, Figures 10 and 11 illustrate DC and AC primal residuals and optimality gaps, respectively.
Figure 10: primal residual (left) and optimality gap (right)
Figure 11: primal residual (left) and optimality gap (right)
Case pglib_opf_case1354_pegase
The dataset pglib_opf_case1354_pegase is part of the PEGASE project, which produced large-scale synthetic European transmission networks intended for stability and optimization studies. This 1354-bus system includes high-density meshed areas characteristic of continental European grids. It is partitioned into eight regions with numerous inter-region connections reflecting the underlying network complexity. The multi-region layout is shown in Fig. 12.
Figure 12: Eight-region topology for case1354_pegase.
Figures 13 and 14 present the DC and AC verification curves for pglib_opf_case1354_pegase.
Figure 13: primal residual (left) and optimality gap (right)
Figure 14: primal residual (left) and optimality gap (right)
Case pglib_opf_case2869_pegase
The system pglib_opf_case2869_pegase represents a large portion of the European transmission network, consisting of 2869 buses, 510 generators, and 4582 branches. The partitioning procedure divides the system into twelve regions, interconnected through ninety-five tie-lines. The resulting multi-region topology is shown in Fig. 15. All buses, generators, and branches are fully accounted for across the regions, and the reference bus (Bus 1314) is assigned to Region 6.
Figure 15: Twelve-region topology for case2869_pegase.
For pglib_opf_case2869_pegase, Figures 16 and 17 illustrate that both DC and AC primal residuals converge to the threshold, while the optimality gaps remain below one percent.
Figure 16: Primal residual (left) and optimality gap (right)
Figure 17: Primal residual (left) and optimality gap (right)
Case pglib_opf_case4661_sdet
The system pglib_opf_case4661_sdet is a large-scale synthetic model developed using the Sustainable Data Evolution Technology (SDET) framework. With 4661 buses and dense connectivity, it reflects realistic large-area transmission dynamics. The system is separated into fourteen regions connected by 284 tie-lines. The partitioned structure is shown in Fig. 18.
Figure 18: Fourteen-region topology for case4661_sdet.
For pglib_opf_case4661_sdet, Figures 19 and 20 indicate that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining below one percent.
Figure 19: primal residual (left) and optimality gap (right)
Figure 20: primal residual (left) and optimality gap (right)
Case pglib_opf_case9241_pegase
The largest benchmark system considered in DPLib is pglib_opf_case9241_pegase, a synthetic European transmission network derived from the PEGASE project. With 9241 buses and more than sixteen thousand branches, it reflects the characteristics of a highly interconnected continental-scale system. The partitioning procedure yields eighteen regions linked by 197 tie-lines. The resulting multi-region topology is shown in Fig. 21.
Figure 21: Eighteen-region topology for case9241_pegase.
For pglib_opf_case9241_pegase, Figures 22 and 23 confirm that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining within one percent, even for this largest benchmark.
Figure 22: primal residual (left) and optimality gap (right)
Figure 23: primal residual (left) and optimality gap (right)